A conducting cylinder of radius 3.25 cm and length 5.6 cm has a total charge of 4.5 x 10^-9 C distributed uniformly on its surface area. Find the potential at a) its surface, and b) 2.5 cm and c) 5.0 cm from the center of the cylinder.

1
Expert's answer
2022-02-23T12:11:53-0500

a)

"d=r,"

"\varphi=2k\lambda\ln\frac dr=0,"

b)

"d<r,"

"\varphi=0,"

c)

"d>r,"

"\varphi=\frac{2kC}{\pi r^2l}\ln\frac dr=190~kV."